OpenStudy (anonymous):

-3c-12=-5+c help! I don't understand

OpenStudy (anonymous):

i dont know .-. I missed the lecture

OpenStudy (anonymous):

first you want to get all the variables on one side and numbers on another: -3c - c = -5 + 12 -4c = 7 then divide both sides by -4 to get the variable by itself c = -7/4

OpenStudy (anonymous):

Okay, your going to want to get c by itself. Hmm, pick the c on the right side of the equation because it's positive. So you have -5+c on the right side of the equation. How do you get the c alone over there?

OpenStudy (anonymous):

-3c-c=-5+12 -4c=7 c=-7/4

OpenStudy (anonymous):

-3c-c=-5+12 -4c=7 divide both sides by -2 and then u'll get, c= -7/4 or 1 3/4 :)

OpenStudy (anonymous):

say you have a, b, c and d as a number and x as a variable: if you have ax + b = cx +d you always want to get all the numbers and variables on one side, so always pick a side where you want variables and numbers: say you want the left to have variables and the right to have numbers: so then you would proceed: ax - cx = d-b then you would divide by that number multiplied by x to get x by itself. if you have x + a = b then you would subtract to get x by itself if you have x - a = b then you would add to get x by itself (add both sides, same as above) if you have ax = b, divide both sides by a to get x by itself and if you have x/a = b then multiply both sides to get x by itself.

OpenStudy (anonymous):

whatever you have on one side with the variable, you basically want to do the opposite to the other side

OpenStudy (anonymous):

Where did you get the 4???